OpenStudy (dls):

Evaluate integration of 1/(1+x^2) from 0 to 1 using Trapezoidal rule taking h=1/4.

OpenStudy (dls):

Attempt : h=1/4 therefore n=4 so according to the formula, \[\LARGE I = h [ \frac{y_0 + y_4}{2} + 2(y_1+y_2+y_3)]\] which turns out to be \[\Large I = \frac{1}{4}(\frac{1}{2}. \frac{18}{17} + 0.5 + 0.2 + 0.1)\] but gives wrong answer.

OpenStudy (triciaal):